# Testing identifying assumptions in Tobit Models

### v2

@@ E.2. Non-Learnability for the IV- Tobit model
As mentioned in Remark 5, the IV-Tobit model is non-verifiable, that is, we cannot learn if the maintained model is the true data generating process based on the sharp equalities proposed in Section 3.2. One can show that by finding an alternative model that is compatible with the equalities (3.4)-(3.12) in all cases in which the IV-Tobit model could not be disregarded. Suppose that,
$\left\{\begin{array}[]{lcl}\tilde{Y}&=&\max(0,\tilde{Y}^{*})\\ \\ \tilde{Y}^{*}&=&G(\beta_{0}+\beta_{1}Z+\tilde{W})\\ \\ \tilde{D}&=&\gamma_{0}+\gamma_{1}Z+\tilde{V}\end{array}\right.$ (E.13)
Where $G$ is the same function as in the previous subsection. Assume that equations 3.4-3.12 hold:
$P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$P(c_{0}\leq Y\leq c_{1},D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$P(c_{0}\leq Y\leq c_{1},D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{W}(c_{0}-\beta_{0}-\beta_{1}z)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$P(c_{1}\leq Y,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)-\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$P(c_{1}\leq Y,D\leq d_{0}|Z=z)$ $=$ $\Phi_{V}(d_{0}-\gamma_{0}-\gamma_{1}z)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$P(c_{1}\leq Y,D\geq d_{1}|Z=z)$ $=$ $1-\Phi_{W}(c_{1}-\beta_{0}-\beta_{1}z)-\Phi_{V}(d_{1}-\gamma_{0}-\gamma_{1}z)$
$+$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
$P(Y=0,d_{0}\leq D\leq d_{1}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$P(Y=0,D\leq d_{0}|Z=z)$ $=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho),$
$P(Y=0,D\geq d_{1}|Z=z)$ $=$ $\Phi_{W}(-\beta_{0}-\beta_{1}z)-\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
Define $\bar{Y}_{z}=G(\beta_{0}+\beta_{1}z+\tilde{W}),\tilde{D}_{z}=\gamma_{0}+\gamma_{1}z+\tilde{V}$ and define the following distribution function:
$P(\bar{Y}_{d}\leq c_{1},\tilde{D}_{z}\leq d_{1},\bar{U}\leq u,\bar{V}\leq v|D=d)=\Phi(\min\{(G^{-1}(c_{1})-\beta_{0}-\beta_{1}z),u\},\min\{d_{1}-\gamma_{0}-\gamma_{1}z,v\};\rho)$
Now note that, for example,
$P(G(c_{0})\leq\tilde{Y}_{z}\leq G(c_{1}),d_{0}\leq\tilde{D}_{z}\leq d_{1})$ $=$ $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$ (E.14)
-- $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{1}-\gamma_{0}-\gamma_{1}z;\rho)$
-- $\Phi_{W,V}(c_{1}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$+$ $\Phi_{W,V}(c_{0}-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$
$=$ $P(c_{0}\leq Y\leq c_{1},d_{0}\leq D\leq d_{1}|Z=z)$
A similar logic holds for the other equalities. But the model is not the IV-tobit model.
Appendix F Simulations for test’s power under violations of normality
Turning our attention to the test’s power under violations of normality of the latent error, we consider the modified DGP in Equation F.7. In this DGP, $F(\cdot)$ denotes an alternative density for the latent error $U$, which we vary across scenarios. Table 6 present the test’s empirical coverage when $F(\cdot)$ is uniform, lognormal and t-student distributions.
$\left\{\begin{array}[]{lcl}Y&=&\max\{Y^{*},0\}\\ Y^{*}&=&D+U\\ D&=&2Z-V\\ \begin{pmatrix}V\\ Z\end{pmatrix}&\sim&\mathcal{N}(\mathbf{0}_{2},\mathbf{\Sigma})\\ \mathbf{\Sigma}&=&\begin{pmatrix}1&0\\ 0&1\\ \end{pmatrix}\\ U\sim F(\cdot)\end{array}\right.$ (F.7)
As expected, the rejection rate for the case of the t-student with 80 degrees of freedom is close to the nominal size of the test, since it represents a very mild violation of the normality assumption. In the other three cases, which deviate significantly from the normality assumption, the test rejects the null hypothesis in most instances. This example indicates that the proposed test is useful in detecting violations of the distributional assumptions in the classic Tobit model.
Table 6. Test power for violations in error structure
$N$ $U$ MSE$(\beta_{0})$ MSE$(\beta_{1})$ $\alpha=10\%$ $\alpha=5\%$ $\alpha=1\%$
5000 t-student (df=80) 0.9987 1.0018 7.21% 4.01% 1.40%
t-student (df=5) 1.027 1.0624 92.20% 81.80% 48.80%
LogNormal 1.3336 0.0896 100.00% 100.00% 100.00%
Uniform 0.892 0.785 100.00% 100.00% 100.00%
8000 t-student (df=80) 0.9999 1.0034 7.00% 4.80% 1.20%
t-student (df=5) 1.0251 1.0653 99.20% 97.80% 87.58%
LogNormal 1.3385 0.0866 100.00% 100.00% 100.00%
Uniform 0.8932 0.7896 100.00% 100.00% 100.00%
10000 t-student (df=80) 1.0016 1.0059 5.42% 2.41% 1.00%
t-student (df=5) 1.0241 1.0642 99.20% 98.40% 90.58%
LogNormal 1.3348 0.0868 100.00% 100.00% 100.00%
Uniform 0.8908 0.7862 100.00% 100.00% 100.00%
Based on $500$ replications.

### v3

@@ E.2. Non-Learnability for the IV- Tobit model
As mentioned in proposition 3, the IV-Tobit model is non-verifiable, that is, we cannot learn if the maintained model is the true data generating process based on the sharp equalities proposed in Section 4.2. One can show that by finding an alternative model that is compatible with the equalities (4.3)-(4.11) in all cases in which the IV-Tobit model could not be disregarded. Suppose that,
$\left\{\begin{array}[]{lcl}\tilde{Y}&=&\max(0,\tilde{Y}^{*})\\ \\ \tilde{Y}^{*}&=&\beta_{0}+\beta_{1}Z+\tilde{W}\\ \\ \tilde{D}&=&\gamma_{0}+\gamma_{1}Z+\tilde{V}\end{array}\right$. (E.15)
$\tilde{W},\tilde{V}$ are random variables that have the next mixed discrete-continuous behavior:
[…]
$f_{\tilde{W}|\tilde{V},Z}(\tilde{w}|\tilde{v},z)=\begin{cases}\Phi(\hat{w}|\tilde{v};\rho)(\frac{e^{z}}{1+e^{z}})\text{ if }\tilde{w}=\hat{w}\\ \Phi(-\beta_{0}-\beta_{1}z|\tilde{v};\rho)-\Phi(\hat{w}|\tilde{v};\rho)(\frac{e^{z}}{1+e^{z}})\text{ if }\tilde{w}=-\beta_{0}-\beta_{1}z\\ \phi(\tilde{w}|\tilde{v};\rho)\text{ if }\tilde{w}\geq-\beta_{0}-\beta_{1}z\\ 0,\text{ otherwise.}\end{cases}$
Where $\hat{w}$ is some real number strictly less than $\beta_{0}-\beta_{1}z$, $\phi(\tilde{w}|\tilde{v},\rho)$ is the normal p.d.f. conditional on $\tilde{v}$ with correlation coefficient $\rho$ and $\Phi(\tilde{w}|\tilde{v};\rho)$ is the normal c.d.f. analog. Note that, $\Phi(\hat{w}|\tilde{v};\rho)(\frac{e^{z}}{1+e^{z}})>0,\phi(\tilde{w}|\tilde{v};\rho)>0$ and $\Phi(-\beta_{0}-\beta_{1}z|\tilde{v};\rho)-\Phi(\hat{w}|\tilde{v};\rho)(\frac{e^{z}}{1+e^{z}})>0$. This distribution assigns probability mass to two points, $\hat{w}$ and $\beta_{0}-\beta_{1}z$ while assigning the conditional normal density above $\beta_{0}-\beta_{1}z$. Thus, above $\beta_{0}-\beta_{1}z$, $\tilde{W}$ behaves like a conditional normal random variable, but below behaves like a discrete random variable with two mass points. Note that, with the previous structure, for values of $\tilde{w}$ less than $\hat{w}$:
[…]
$=$ $\int_{-\infty}^{\tilde{v}}P(\tilde{W}\leq\hat{w}|\tilde{V}=v)\phi(v)dv\frac{e^{z}}{1+e^{z}}$
Furthermore, note that:
$\int_{-\infty}^{\tilde{v}}\Phi(\hat{w}|\tilde{v};\rho)\phi(\tilde{v})d\tilde{v}$ $=$ $\int_{-\infty}^{\tilde{w}}\int_{-\infty}^{\tilde{v}}\phi(\hat{w}|\tilde{v};\rho)\phi(\tilde{v})d\tilde{v}d\tilde{w}=\int_{-\infty}^{\tilde{w}}\int_{-\infty}^{\tilde{v}}\phi(\hat{w},\tilde{v};\rho)d\tilde{v}d\tilde{w}$
$=$ $\Phi(\tilde{w},\tilde{v};\rho)$
With this, we can then say that:
$F_{\tilde{W},\tilde{V}|Z}(\tilde{w},\tilde{v}|z)=\begin{cases}\Phi(\hat{w},\tilde{v};\rho)\frac{e^{z}}{1+e^{z}}\text{ if }\tilde{w}=\hat{w},-\infty<\tilde{v}\\ \Phi(-\beta_{0}-\beta_{1}z,\tilde{v};\rho)\text{ if }\hat{w}<\tilde{w}\leq-\beta_{0}-\beta_{1}z,-\infty<\tilde{v}\\ \Phi(\tilde{w},\tilde{v};\rho)\text{ if }\tilde{w}>-\beta_{0}+\beta_{1}z,-\infty<\tilde{v}\\ 0,\text{ otherwise}\end{cases}$
This joint c.d.f., which integrates to $1$ when $\tilde{w}$ and $\tilde{v}$ goes to $\infty$ is associated with the following mixed discrete-continuous ”density”:
$f_{\tilde{W},\tilde{V}|Z}(\tilde{w},\tilde{v}|z)=\begin{cases}\Phi(\hat{w},\tilde{v};\rho)\frac{e^{z}}{1+e^{z}}\text{ if }\tilde{w}=\hat{w},\forall\tilde{v}\\ \Phi(-\beta_{0}-\beta_{1}z,\tilde{v};\rho)-\Phi(\hat{w},\tilde{v};\rho)\frac{e^{z}}{1+e^{z}}\text{ if }\tilde{w}=-\beta_{0}-\beta_{1}z,\forall\tilde{v}\\ \phi(\tilde{w},\tilde{v};\rho)\text{ if }\tilde{w}>-\beta_{0}+\beta_{1}z,\forall\tilde{v}\end{cases}$
Assume that equations 4.3-4.11 hold. Then, define $\bar{Y}_{z}=\beta_{0}+\beta_{1}z+\tilde{W},\tilde{D}_{z}=\gamma_{0}+\gamma_{1}z+\tilde{V}$
$P(\bar{Y}_{z}<0,D_{z}\leq d_{0})$ $=$ $P(\tilde{W}\leq-\beta_{0}-\beta_{1}z,\tilde{V}\leq d_{0}-\gamma_{0}-\gamma_{1}z)$ (E.16)
$=$ $\Phi_{W,V}(-\beta_{0}-\beta_{1}z,d_{0}-\gamma_{0}-\gamma_{1}z;\rho)$ (E.17)
$=$ $P(Y=0,D\leq d_{0}|Z=z)$ (E.18)
The last equality is due to the testable equalities holding. A similar display holds for any $c_{0},c_{1}>0,d_{1},d_{0}$. Thus a similar logic holds for the other equalities. So this is an example of a model that satisfies all the equalities, and is observationally equivalent to the IV-Tobit, even though the underlying latent error distributions does not satisfy the conditions for the IV-Tobit. Hence, we cannot falsify the assumed IV-Tobit model.